# Adaptive Prompt: Unlocking the Power of Visual Prompt Tuning

### v1

@@ A.1.1 Local part
for any $1\leq\tau\leq 6$ and $1\leq j\leq L$. Here, from the definition of $m_{n}$, at least one coefficient among $\{\lambda_{0,j},\lambda_{1,j},\lambda_{2,j},\lambda_{3,j}\}_{j:|\mathcal{V}_{j}|=1}$, $\{\lambda_{0,j},\lambda_{1,j},\lambda_{2,j},\lambda_{3,j},\lambda_{4,j},\lambda_{5,j},\lambda_{6,j}\}_{j:|\mathcal{V}_{j}|>1}$ is different from 0. Then, the equation $\liminf_{n\to\infty}\dfrac{\left|f_{G_{n}}({\bm{X}})-f_{G_{*}}({\bm{X}})\right|}{m_{n}\mathcal{D}_{1n}}=0$, or equivalently, $\liminf_{n\to\infty}\dfrac{\left|Q_{n}({\bm{X}})\right|}{m_{n}\mathcal{D}_{1n}}=0$ leads to
@@
for almost surely ${\bm{X}}$. However, the new equation implies that all the coefficients $\{\lambda_{0,j},\lambda_{1,j},\lambda_{2,j},\lambda_{3,j}\}_{j:|\mathcal{V}_{j}|=1}$, $\{\lambda_{0,j},\lambda_{1,j},\lambda_{2,j},\lambda_{3,j},\lambda_{4,j},\lambda_{5,j},\lambda_{6,j}\}_{j:|\mathcal{V}_{j}|>1}$ are 0, which is a contradiction.
As a consequence, we obtain
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{G}_{L^{\prime}}(\Theta):\mathcal{D}_{1}(G,G_{*})\leq\varepsilon}\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{1}(G,G_{*})>0.$
A.1.2 Global part
From the local part, there exists a positive constant $\varepsilon^{\prime}$ such that
@@ A.1.2 Global part
Therefore, it is sufficient to prove that
@@
Assume by contrary, then we can find a sequence of mixing measures $G^{\prime}_{n}:=\sum_{j^{\prime}=1}^{L^{\prime}}\exp(b_{n,j^{\prime}})\delta_{(W_{n,1j^{\prime}},W_{n,2})}$ in $\mathcal{G}_{L^{\prime}}(\Theta)$ such that as $n\to\infty$, we have
@@
which indicates that $\|f_{G^{\prime}_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}\to 0$ as $n\to\infty$.
Note that $\Theta$ is a compact set, there exists a mixing measure $G^{\prime}$ in $\mathcal{G}_{L^{\prime}}(\Theta)$ such that one of $G^{\prime}_{n}$’s subsequences converges to $G^{\prime}$. Since $\mathcal{D}_{1}(G^{\prime}_{n},G_{*})>\varepsilon^{\prime}$, we deduce that $\mathcal{D}_{1}(G^{\prime},G_{*})>\varepsilon^{\prime}$. By invoking the Fatou’s lemma, we have that
@@
Thus, we have $f_{G^{\prime}}=f_{G_{*}}$ for $\mu-$almost surely ${\bm{X}}$. From the identifiability property, we deduce that $G^{\prime}\equiv G_{*}$. It follows that $\mathcal{D}_{1}(G^{\prime},G_{*})=0$, contradicting the fact that $\mathcal{D}_{1}(G^{\prime},G_{*})>\varepsilon^{\prime}>0$. Hence, the proof is completed.
Identifiability property.
We now show that if $f_{G}({\bm{X}})=f_{G_{*}}({\bm{X}})$ for almost every ${\bm{X}}$, then it follows that $G\equiv G_{*}$.
For any $G\in\mathcal{G}_{L^{\prime}}(\Xi)$ and $G_{*}$, let us denote
$\mathrm{softmax}_{G}(u)$ $=\dfrac{\exp(u)}{\sum_{k=1}^{N}\exp({\bm{X}}^{\top}A^{0}_{k}{\bm{X}}+a^{0}_{k})+\sum_{j^{\prime}=1}^{L^{\prime}}\exp((BW_{2}\sigma(W_{1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{j^{\prime}})},$
$\mathrm{softmax}_{G_{*}}(u_{*})$ $=\dfrac{\exp(u_{*})}{\sum_{k=1}^{N}\exp({\bm{X}}^{\top}A^{0}_{k}{\bm{X}}+a^{0}_{k})+\sum_{j^{\prime}=1}^{L}\exp((BW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{*,j^{\prime}})},$
where
$u$$\in\{{\bm{X}}^{\top}A^{0}_{j}{\bm{X}}+a^{0}_{j};(BW_{2}\sigma(W_{1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{j^{\prime}}:j\in[N],j^{\prime}\in[L^{\prime}]\},$
$u_{*}$$\in\{{\bm{X}}^{\top}A^{0}_{j}{\bm{X}}+a^{0}_{j};(BW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{*,j^{\prime}}:j\in[N],j^{\prime}\in[L]\}.$
Since $f_{G}({\bm{X}})=f_{G_{*}}({\bm{X}})$ for almost every ${\bm{X}}$, we have
$\sum_{j=1}^{N}\mathrm{softmax}_{G}({\bm{X}}^{\top}A^{0}_{j}{\bm{X}}+a^{0}_{j}))h({\bm{X}},\eta^{0}_{j})+\sum_{j^{\prime}=1}^{L^{\prime}}\mathrm{softmax}_{G}((BW_{2}\sigma(W_{1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{j^{\prime}})CW_{2}\sigma(W_{1j^{\prime}}{\bm{X}})$
$=\sum_{j=1}^{N}\mathrm{softmax}_{G_{*}}({\bm{X}}^{\top}A^{0}_{j}{\bm{X}}+a^{0}_{j}))h({\bm{X}},\eta^{0}_{j})+\sum_{j^{\prime}=1}^{L}\mathrm{softmax}_{G_{*}}((BW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{*,j^{\prime}})CW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}).$ (27)
Thus, we must have that $L=L^{\prime}$. As a result, we obtain that
$\{\mathrm{softmax}_{G}((BW_{2}\sigma(W_{1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{j^{\prime}}):j^{\prime}\in[L]\}$ =$=\{\mathrm{softmax}_{G_{*}}((BW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{*,j^{\prime}}):j^{\prime}\in[L^{\prime}]\},$
for almost every ${\bm{X}}$. We may assume that
$\mathrm{softmax}_{G}((BW_{2}\sigma(W_{1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{j^{\prime}})=\mathrm{softmax}_{G_{*}}((BW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{*,j^{\prime}}),$
for any $j^{\prime}\in[L]$, for almost every ${\bm{X}}$. Since the softmax function is invariant to translation, this result indicates that $b_{j^{\prime}}=b_{*,j^{\prime}}+r$ for some $r\in\mathbb{R}$ and for any $j^{\prime}\in[L]$. Then, the equation (27) can be reduced to
@@
for almost every ${\bm{X}}$. Next, we will partition the index set $[L]$ into $m$ subsets $K_{1},K_{2},\ldots,K_{m}$ where $m\leq L$, such that $\exp{(b_{j})}=\exp{(b_{*,j^{\prime}})}$ for any $j,j^{\prime}\in K_{i}$ and $i\in[m]$. It follows that $\exp{(b_{j})}\neq\exp{(b_{*,j^{\prime}})}$ when $j,j^{\prime}$ do not belong to the same set $K_{i}$. Thus, we can rewrite equation (28) as
$\sum_{i=1}^{m}\sum_{j\in K_{i}}\exp{(b_{j})}\exp{((BW_{2}\sigma(W_{1j}{\bm{X}}))^{\top}{\bm{X}})}CW_{2}\sigma(W_{1j}{\bm{X}})$
$=\sum_{i=1}^{m}\sum_{j^{\prime}\in K_{i}}\exp{(b_{*,j})}\exp{((BW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}})}CW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}),$
for almost surely ${\bm{X}}$. Given the above equation, for each $i\in[m]$, we obtain that
$\{W_{2}\sigma(W_{1j}{\bm{X}}):j\in K_{i}\}=\{W_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}):j^{\prime}\in K_{i}\},$
Since the activation function $\sigma$ is identifiable, the above result indicates that
$\{(W_{1j},W_{2}):j\in K_{i}\}=\{(W_{*,1j^{\prime}},W_{2}):j^{\prime}\in K_{i}\},$
Consequently, we get that
$\sum_{i=1}^{m}\sum_{j\in K_{i}}\exp{(b_{j})}\delta_{(W_{1j},W_{2})}=\sum_{i=1}^{m}\sum_{j^{\prime}\in K_{i}}\exp{(b_{*,j})}\delta_{(W_{*,1j^{\prime}},W_{*,2})},$
which implies that $G\equiv G_{*}$. As a consequence, the proof is completed.
@@ Appendix B Additional Theoretical Results for Visual Adaptive Prompt Tuning
Setting. We assume that the i.i.d. samples of size $n$: $({\bm{X}}_{1},Y_{1}),({\bm{X}}_{2},Y_{2}),\ldots,({\bm{X}}_{n},Y_{n})\in\mathbb{R}^{d}\times\mathbb{R}^{d^{\prime}}$ are generated from the model:
$Y_{i}=f_{\tilde{G}_{*}}({\bm{X}}_{i})+\varepsilon_{i},\quad i=1,2,\ldots,n,$ (29)
where $\varepsilon_{1},\ldots,\varepsilon_{n}$ are independent Gaussian noise variables such that $\mathbb{E}[{\varepsilon_{i}}|{\bm{X}}_{i}]=0$ and $\mathrm{Var}(\varepsilon_{i}|{\bm{X}}_{i})=\nu^{2}I_{d^{\prime}}$ for all $1\leq i\leq n$. Additionally, we assume that ${\bm{X}}_{1},{\bm{X}}_{2},\ldots,{\bm{X}}_{n}$ are i.i.d. samples from some probability distribution $\mu$. The regression function $f_{\tilde{G}_{*}}(\cdot)$ in equation (29) then takes the form of a prefix MoE model with $N$ pre-trained experts and $L$ unknown experts,
$f_{\tilde{G}_{*}}({\bm{X}}):=\sum_{j=1}^{N}\frac{\exp({\bm{X}}^{\top}A^{0}_{j}{\bm{X}}+a^{0}_{j})}{\tilde{D}_{f}({\bm{X}})}\cdot h({\bm{X}},\eta^{0}_{j})+\sum_{j^{\prime}=1}^{L}\frac{\exp((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{*,j^{\prime}})}{\tilde{D}_{f}({\bm{X}})}\cdot CW_{*,2}W_{*,1j^{\prime}}{\bm{X}},$ (30)
where $\tilde{D}_{f}({\bm{X}}):=\sum_{k=1}^{N}\exp({\bm{X}}^{\top}A^{0}_{k}{\bm{X}}+a^{0}_{k})+\sum_{j^{\prime}=1}^{L}\exp((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{*,j^{\prime}})$, while $\tilde{G}_{*}:=\sum_{j^{\prime}=1}^{L}\exp(b_{*,j^{\prime}})\delta_{W_{*,2}W_{*,1j^{\prime}}}$ denotes a mixing measure, i.e., a weighted sum of Dirac measures $\delta$, associated with unknown parameters $(b_{*,j^{\prime}},W_{*,2}W_{*,1j^{\prime}})_{j^{\prime}=1}^{L}$ in the parameter space $\Theta\subset\mathbb{R}\times\mathbb{R}^{d\times d}$. At the same time, the values of the matrix $A^{0}_{j}$, the expert parameter $\eta^{0}_{j}$, and the bias parameter $a^{0}_{j}$ are known for all $1\leq j\leq N$. Additionally, the matrices $B\in\mathbb{R}^{d\times d}$ and $C\in\mathbb{R}^{d^{\prime}\times d}$ are given.
@@
$\tilde{G}_{n}:=\operatorname*{arg\,min}_{G\in\mathcal{G}_{L^{\prime}}(\Theta)}\sum_{i=1}^{n}\Big{(}Y_{i}-f_{G}({\bm{X}}_{i})\Big{)}^{2},$ (31)
@@
$\mathcal{D}_{2}({G},\tilde{G}_{*}):=\sum_{j^{\prime}=1}^{L}\Big{|}\sum_{i\in\mathcal{V}_{j^{\prime}}}\exp(b_{i})-\exp(b_{*,j^{\prime}})\Big{|}$ $+\sum_{j^{\prime}\in[L]:|\mathcal{V}_{j^{\prime}}|=1}\sum_{i\in\mathcal{V}_{j^{\prime}}}\exp(b_{i})\|\Delta W_{2}W_{1ij^{\prime}}\|$
@@
Given the least square estimator $\tilde{G}_{n}$ defined in equation (31), we have that
$\mathcal{D}_{2}(\tilde{G}_{n},\tilde{G}_{*})=\mathcal{O}_{P}(\sqrt{\log(n)/n}).$
@@
Given the convergence rate of regression function estimation in Proposition 1, our goal is to demonstrate the following inequality:
$\inf_{G\in\mathcal{G}_{L^{\prime}}(\Theta)}\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{2}(G,G_{*})>0.$
We divide the proof of the above inequality into local and global parts.
Local part:
We will demonstrate that
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{G}_{L^{\prime}}(\Theta):\mathcal{D}_{2}(G,\tilde{G}_{*})\leq\varepsilon}\|f_{G}-f_{\tilde{G}_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{2}(G,\tilde{G}_{*})>0.$
Assume by contrary that the above claim does not hold. Then, there exists a sequence of mixing measures $G_{n}:=\sum_{j^{\prime}=1}^{L^{\prime}}\exp(b_{n,j^{\prime}})\delta_{W_{n,2}W_{n,1j^{\prime}}}$ in $\mathcal{G}_{L^{\prime}}(\Theta)$ such that as $n\to\infty$, we have
$\left\{\begin{matrix}\mathcal{D}_{2n}:=\mathcal{D}_{2}(G_{n},\tilde{G}_{*})\to 0,\\ \|f_{G_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{2n}\to 0.\end{matrix}\right.$
To ease the ensuing presentation, we also denote $\mathcal{V}_{j}^{n}:=\mathcal{V}_{j}(G_{n})$ as a Voronoi cell of $G_{n}$ generated by the $j$-th components of $\tilde{G}_{*}$. Since our arguments are asymptotic, we may assume that those Voronoi cells do not depend on the sample size, i.e., $\mathcal{V}_{j}=\mathcal{V}_{j}^{n}$. Therefore, we can represent the Voronoi loss $\mathcal{D}_{2n}$ as follows:
@@
Additionally, since $\mathcal{D}_{2n}\to 0$ as $n\to\infty$, we have $\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})\to\exp(b_{*,j})$, $W_{n,2}W_{n,1i}\to W_{*,2}W_{*,1j^{\prime}}$ for any $i\in\mathcal{V}_{j},j\in[L]$. Throughout this proof, we assume WLOG that $B=Id_{d}$, $C=Id_{d}$ and $r=1$ for simplicity with a note that our techniques can totally be generalized to the general case. Now, we divide the proof of the local part into three steps as follows:

### v3

@@ A.1 Proof of Theorem 1
for any $1\leq\tau\leq 6$ and $1\leq j\leq L$. By the definition of $m_{n}$, at least one coefficient in the set $\{\lambda_{0,j},\lambda_{1,j},\lambda_{2,j},\lambda_{3,j}\}_{j:|\mathcal{V}_{j}|=1}$, $\{\lambda_{0,j},\lambda_{1,j},\lambda_{2,j},\lambda_{3,j},\lambda_{4,j},\lambda_{5,j},\lambda_{6,j}\}_{j:|\mathcal{V}_{j}|>1}$ must be nonzero. The limit $\liminf_{n\to\infty}\dfrac{\left|f_{G_{n}}({\bm{X}})-f_{G_{*}}({\bm{X}})\right|}{m_{n}\mathcal{D}_{1n}}=0$, or equivalently, $\liminf_{n\to\infty}\dfrac{\left|Q_{n}({\bm{X}})\right|}{m_{n}\mathcal{D}_{1n}}=0$ leads to
@@
for almost surely ${\bm{X}}$. However, that equation implies that all the coefficients $\{\lambda_{0,j},\lambda_{1,j},\lambda_{2,j},\lambda_{3,j}\}_{j:|\mathcal{V}_{j}|=1}$, $\{\lambda_{0,j},\lambda_{1,j},\lambda_{2,j},\lambda_{3,j},\lambda_{4,j},\lambda_{5,j},\lambda_{6,j}\}_{j:|\mathcal{V}_{j}|>1}$ are 0. It is a contradiction to the hypothesis that at least one coefficient among these coefficients is different from 0.
Consequently, we deduce that
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{G}_{L^{\prime}}(\Theta):\mathcal{D}_{1}(G,G_{*})\leq\varepsilon}\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{1}(G,G_{*})>0,$
which suggests that the conclusion of the local part of the inequality (20) is proved.
A.1.2 Global part.
From the result of the local part of the inequality (20), we can find a positive constant $\varepsilon^{\prime}$ such that the following inequality holds:
@@
To obtain the conclusion of the theorem, we only need to demonstrate that
@@
We prove the claim by contradiction. Indeed, by assuming the claim does not hold, it implies that there exists a sequence of $G^{\prime}_{n}:=\sum_{j^{\prime}=1}^{L^{\prime}}\exp(b_{n,j^{\prime}})\delta_{(W_{n,1j^{\prime}},W_{n,2})}$ in the set $\mathcal{G}_{L^{\prime}}(\Theta)$ such that
@@
as long as $n$ approaches the infinity. These results prove that $\|f_{G^{\prime}_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}\to 0$ as $n$ goes to infinity.
From the hypothesis, the parameter space $\Theta$ is compact. Therefore, one of $G^{\prime}_{n}$’s subsequences converges to some mixing measure $G^{\prime}$ where $G^{\prime}$ lies in the space $\mathcal{G}_{L^{\prime}}(\Theta)$. From the hypothesis, we have $\mathcal{D}_{1}(G^{\prime}_{n},G_{*})>\varepsilon^{\prime}$. By taking the limit of both sides as $n\to\infty$, we obtain that $\mathcal{D}_{1}(G^{\prime},G_{*})\geq\varepsilon^{\prime}$.
An application of the Fatou’s lemma leads to the following result:
@@
This inequality is only possible if $f_{G^{\prime}}=f_{G_{*}}$ for almost surely ${\bm{X}}$.
According to the identifiability of the function $f_{G}({\bm{X}})$, that equation only holds when $G^{\prime}\equiv G_{*}$. As a consequence, we obtain that $\mathcal{D}_{1}(G^{\prime},G_{*})=0$. It contradicts to the assumption that $\mathcal{D}_{1}(G^{\prime},G_{*})\geq\varepsilon^{\prime}>0$. Hence, the proof of the global part is completed. We achieve the conclusion of the theorem.
Proof for the identifiability property.
The key claim that we aim to show is that if the equation $f_{G}({\bm{X}})=f_{G_{*}}({\bm{X}})$ holds for almost surely ${\bm{X}}$, we achieve that $G\equiv G_{*}$, namely, the two mixing measures are identical.
From the hypothesis, as $f_{G}({\bm{X}})=f_{G_{*}}({\bm{X}})$ for almost all ${\bm{X}}$, we achieve that
$\sum_{j=1}^{N}\frac{\exp({\bm{X}}^{\top}A^{0}_{j}{\bm{X}}+a^{0}_{j}))}{D_{f,G}({\bm{X}})}h({\bm{X}},\eta^{0}_{j})+\sum_{j^{\prime}=1}^{\tilde{L}}\frac{\exp((BW_{2}\sigma(W_{1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{j^{\prime}})}{D_{f,G}({\bm{X}})}CW_{2}\sigma(W_{1j^{\prime}}{\bm{X}})$
$=\sum_{j=1}^{N}\frac{\exp({\bm{X}}^{\top}A^{0}_{j}{\bm{X}}+a^{0}_{j}))}{D_{f,G_{*}}({\bm{X}})}h({\bm{X}},\eta^{0}_{j})+\sum_{j^{\prime}=1}^{L}\frac{\exp((BW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{*,j^{\prime}})}{D_{f,G_{*}}({\bm{X}})}CW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}),$ (27)
where $G=\sum_{j=1}^{\tilde{L}}\exp(b_{j})\delta_{(W_{1j},W_{2})}$. Furthermore, we define
$D_{f,G_{*}}({\bm{X}})$ $=\sum_{k=1}^{N}\exp({\bm{X}}^{\top}A^{0}_{k}{\bm{X}}+a^{0}_{k})+\sum_{j^{\prime}=1}^{L}\exp((BW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{*,j^{\prime}}),$
$D_{f,G}({\bm{X}})$ $=\sum_{k=1}^{N}\exp({\bm{X}}^{\top}A^{0}_{k}{\bm{X}}+a^{0}_{k})+\sum_{j^{\prime}=1}^{\tilde{L}}\exp((BW_{2}\sigma(W_{1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{j^{\prime}}).$
That equation implies that the number of atoms of $G$ and $G_{*}$ should be identical, namely, we have $L=\tilde{L}$. Therefore, the following result holds
$\Bigg{\{}\frac{\exp((BW_{2}\sigma(W_{1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{j^{\prime}})}{D_{f,G}({\bm{X}})}:j^{\prime}\in[L]\Bigg{\}}$ =$=\Bigg{\{}\frac{\exp((BW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{*,j^{\prime}})}{D_{f,G_{*}}({\bm{X}})}:j^{\prime}\in[L]\Bigg{\}},$
for almost surely ${\bm{X}}$. By relabelling the indices of these two sets, we can assume without loss of generality that
$\frac{\exp((BW_{2}\sigma(W_{1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{j^{\prime}})}{D_{f,G}({\bm{X}})}$ $=\frac{\exp((BW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{*,j^{\prime}})}{D_{f,G_{*}}({\bm{X}})},$
for any index $j^{\prime}\in[L]$ and for almost surely ${\bm{X}}$. From the invariance to translation property of the softmax function, the equation (27) becomes
@@
for almost surely ${\bm{X}}$. This equation suggests that there exists a partition $K_{1},K_{2},\ldots,K_{m}$ of the set $[L]$ for some $m$ such that we have $\exp(b_{j_{1}})=\exp(b_{*,j_{2}})$ for any $j_{1},j_{2}\in K_{i}$ and for any $i\in[m]$. According to that result, the equation (27) can be rewritten as follows:
$\sum_{i=1}^{m}\sum_{j_{1}\in K_{i}}\exp{(b_{j_{1}})}\exp{((BW_{2}\sigma(W_{1j_{1}}{\bm{X}}))^{\top}{\bm{X}})}CW_{2}\sigma(W_{1j_{1}}{\bm{X}})$
$=\sum_{i=1}^{m}\sum_{j_{2}\in K_{i}}\exp{(b_{*,j_{2}})}\exp{((BW_{*,2}\sigma(W_{*,1j_{2}}{\bm{X}}))^{\top}{\bm{X}})}CW_{*,2}\sigma(W_{*,1j_{2}}{\bm{X}}),$
for almost surely ${\bm{X}}$. That equation proves that
$\{W_{2}\sigma(W_{1j_{1}}{\bm{X}}):j_{1}\in K_{i}\}=\{W_{*,2}\sigma(W_{*,1j_{2}}{\bm{X}}):j_{2}\in K_{i}\},$
for any $i\in[m]$. Since the activation function $\sigma$ is identifiable, that result indicates that
$\{(W_{1j_{1}},W_{2}):j_{1}\in K_{i}\}=\{(W_{*,1j_{2}},W_{2}):j_{2}\in K_{i}\}.$
As a consequence, we arrive at the following result:
$\sum_{i=1}^{m}\sum_{j_{1}\in K_{i}}\exp{(b_{j_{1}})}\delta_{(W_{1j_{1}},W_{2})}=\sum_{i=1}^{m}\sum_{j_{2}\in K_{i}}\exp{(b_{*,j_{2}})}\delta_{(W_{*,1j_{2}},W_{*,2})}.$
It is equivalent to $G\equiv G_{*}$. As a consequence, we obtain the conclusion of the identifiability property of the function $f_{G}$.
@@ Appendix B Additional Theoretical Results for Visual Adaptive Prompt Tuning
Problem setup. We assume that the i.i.d. samples of size $n$: $({\bm{X}}_{1},Y_{1}),({\bm{X}}_{2},Y_{2}),\ldots,({\bm{X}}_{n},Y_{n})\in\mathbb{R}^{d}\times\mathbb{R}^{d^{\prime}}$ are generated from the model:
$Y_{i}=f_{\widetilde{G}_{*}}({\bm{X}}_{i})+\varepsilon_{i},\quad i=1,2,\ldots,n,$ (29)
where $\varepsilon_{1},\ldots,\varepsilon_{n}$ are independent Gaussian noise variables such that $\mathbb{E}[{\varepsilon_{i}}|{\bm{X}}_{i}]=0$ and $\mathrm{Var}(\varepsilon_{i}|{\bm{X}}_{i})=\nu^{2}I_{d^{\prime}}$ for all $1\leq i\leq n$. Additionally, we assume that ${\bm{X}}_{1},{\bm{X}}_{2},\ldots,{\bm{X}}_{n}$ are i.i.d. samples from some probability distribution $\mu$. The regression function $f_{\widetilde{G}_{*}}(\cdot)$ in equation (29) then takes the form of a prefix MoE model with $N$ pre-trained experts and $L$ unknown experts,
$f_{\widetilde{G}_{*}}({\bm{X}}):=\sum_{j=1}^{N}$ $\frac{\exp({\bm{X}}^{\top}A^{0}_{j}{\bm{X}}+a^{0}_{j})}{\tilde{D}_{f}({\bm{X}})}\cdot h({\bm{X}},\eta^{0}_{j})$
$+\sum_{j^{\prime}=1}^{L}\frac{\exp((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{*,j^{\prime}})}{\tilde{D}_{f}({\bm{X}})}\cdot CW_{*,2}W_{*,1j^{\prime}}{\bm{X}},$ (30)
where $\tilde{D}_{f,\widetilde{G}_{*}}({\bm{X}}):=\sum_{k=1}^{N}\exp({\bm{X}}^{\top}A^{0}_{k}{\bm{X}}+a^{0}_{k})+\sum_{j^{\prime}=1}^{L}\exp((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{*,j^{\prime}})$, while $\widetilde{G}_{*}:=\sum_{j^{\prime}=1}^{L}\exp(b_{*,j^{\prime}})\delta_{W_{*,2}W_{*,1j^{\prime}}}$ denotes a mixing measure, i.e., a weighted sum of Dirac measures $\delta$, associated with unknown parameters $(b_{*,j^{\prime}},W_{*,2}W_{*,1j^{\prime}})_{j^{\prime}=1}^{L}$ in the parameter space $\Theta\subset\mathbb{R}\times\mathbb{R}^{d\times d}$. At the same time, the values of the matrix $A^{0}_{j}$, the expert parameter $\eta^{0}_{j}$, and the bias parameter $a^{0}_{j}$ are known for all $1\leq j\leq N$. Additionally, the matrices $B\in\mathbb{R}^{d\times d}$ and $C\in\mathbb{R}^{d^{\prime}\times d}$ are given.
@@
$\widetilde{G}_{n}:=\operatorname*{arg\,min}_{G\in\mathcal{G}_{L^{\prime}}(\Theta)}\sum_{i=1}^{n}\Big{(}Y_{i}-f_{G}({\bm{X}}_{i})\Big{)}^{2},$ (31)
@@
$\mathcal{D}_{2}({G},\widetilde{G}_{*}):=\sum_{j^{\prime}=1}^{L}\Big{|}\sum_{i\in\mathcal{V}_{j^{\prime}}}\exp(b_{i})-\exp(b_{*,j^{\prime}})\Big{|}$ $+\sum_{j^{\prime}\in[L]:|\mathcal{V}_{j^{\prime}}|=1}\sum_{i\in\mathcal{V}_{j^{\prime}}}\exp(b_{i})\|\Delta W_{2}W_{1ij^{\prime}}\|$
@@
Given the least square estimator $\widetilde{G}_{n}$ defined in equation (31), we have that
$\mathcal{D}_{2}(\widetilde{G}_{n},\widetilde{G}_{*})=\mathcal{O}_{P}([\log(n)/n]^{\frac{1}{2}}).$
@@
Based on the convergence rate of regression function estimation presented in Proposition 1, our objective is to establish the following inequality:
$\inf_{G\in\mathcal{G}_{L^{\prime}}(\Theta)}\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{2}(G,G_{*})>0.$ (32)
We partition the proof of the above inequality into local and global parts.
B.0.1 Local part.
The local part of the inequality (32) corresponds to the following inequality:
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{G}_{L^{\prime}}(\Theta):\mathcal{D}_{2}(G,\widetilde{G}_{*})\leq\varepsilon}\frac{\|f_{G}-f_{\widetilde{G}_{*}}\|_{L_{2}(\mu)}}{\mathcal{D}_{2}(G,\widetilde{G}_{*})}>0.$
We assume that the above inequality does not hold. Then, we can find a sequence of measures $G_{n}:=\sum_{j^{\prime}=1}^{L^{\prime}}\exp(b_{n,j^{\prime}})\delta_{W_{n,2}W_{n,1j^{\prime}}}$ in $\mathcal{G}_{L^{\prime}}(\Theta)$ such that as $n\to\infty$, we have
$\left\{\begin{matrix}\mathcal{D}_{2n}:=\mathcal{D}_{2}(G_{n},\widetilde{G}_{*})\to 0,\\ \|f_{G_{n}}-f_{\widetilde{G}_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{2n}\to 0.\end{matrix}\right.$
Similar to the proof of Theorem 1, for the sake of the presentation, $\mathcal{V}_{j}^{n}:=\mathcal{V}_{j}(G_{n})$ is denoted as a Voronoi cell of $G_{n}$ generated by the $j$-th components of the true measure $\widetilde{G}_{*}$. Given that the ensuing arguments are asymptotic, without loss of generality we assume that those Voronoi cells do not depend on the sample size, i.e., we have $\mathcal{V}_{j}=\mathcal{V}_{j}^{n}$ for all $n$ and $1\leq j\leq L$. Therefore, the Voronoi loss $\mathcal{D}_{2n}$ can be rewritten as follows:
@@
From the hypothesis, since $\mathcal{D}_{2n}\to 0$ as $n$ approaches $\infty$, we find that $\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})\to\exp(b_{*,j})$ and $W_{n,2}W_{n,1i}\to W_{*,2}W_{*,1j^{\prime}}$ for any $i\in\mathcal{V}_{j}$ and $j\in[L]$. Throughout this proof, we assume WLOG that $B=Id_{d}$, $C=Id_{d}$ and $r=1$ for simplicity with a note that our techniques can totally be generalized to the general case. We partition the proof of the local part into three steps, as detailed below: